# Pariksha : A Large-Scale Investigation of Human-LLM Evaluator Agreement on Multilingual and Multi-Cultural Data

### v1

@@ Abstract
Evaluation of multilingual Large Language Models (LLMs) is challenging due to a variety of factors -- the lack of benchmarks with sufficient linguistic diversity, contamination of popular benchmarks into LLM pre-training data and the lack of local, cultural nuances in translated benchmarks. In this work, we study human and LLM-based evaluation in a multilingual, multi-cultural setting. We evaluate 30 models across 10 Indic languages by conducting 90K human evaluations and 30K LLM-based evaluations and find that models such as GPT-4o and Llama-3 70B consistently perform best for most Indic languages. We build leaderboards for two evaluation settings - pairwise comparison and direct assessment and analyse the agreement between humans and LLMs. We find that humans and LLMs agree fairly well in the pairwise setting but the agreement drops for direct assessment evaluation especially for languages such as Bengali and Odia. We also check for various biases in human and LLM-based evaluation and find evidence of self-bias in the GPT-based evaluator. Our work presents a significant step towards scaling up multilingual evaluation of LLMs. [^1] We plan to release all evaluation artifacts post acceptance
@@
Pariksha : A Large-Scale Investigation of Human-LLM Evaluator Agreement on Multilingual and Multi-Cultural Data
@@ 1 Introduction
Figure 1: Evaluation pipeline (1) We curate a diverse set of evaluation prompts with the help of native speakers. (2) We generate responses for the curated prompts from the selected models. (3) We evaluate generated responses in two settings (direct assessment and pairwise comparison) by both Humans and an LLM. (4) We construct leaderboards using scores obtained and analyze the agreement between human and LLM evaluators.
@@
However, standard benchmarking suffers from the following issues: many popular benchmarks are available on the web and have already been consumed in the training data of LLMs, rendering them unsuitable for fair evaluation. This phenomenon is known as test dataset contamination, and recent work Ravaut et al. (2024); Golchin and Surdeanu (2024); Dong et al. (2024); Oren et al. (2024); Deng et al. (2024) has suggested that contamination can occur not only during pre-training, but also during fine-tuning and evaluation Balloccu et al. (2024). This calls for dynamic benchmarking with the help of humans Chiang et al. (2024); Yang et al. (2023). Although, human evaluation is considered the gold standard, it can be expensive and time consuming. Due to this, the use of LLM-evaluators, where an LLM itself is used to evaluate the output of another LLM (sometimes the same LLM) has become very popular.
@@
Most studies on LLM training and evaluation focus on English. Recent work has shown that LLMs perform worse on non-English languages, particularly those written in scripts other than the Latin script, and under-resourced languages Ahuja et al. (2023, 2024); Asai et al. (2024). Studies on cultural values in LLMs have also shown that frontier models such as GPT-4 align more closely to Western, Rich, Industrialized norms Rao et al. (2023). This has led to a proliferation of models being built for specific languages, cultures and regions such as Indic, Arabic, African, Chinese, European, and Indonesian (Gala et al., 2024; Sengupta et al., 2023; Zeng et al., 2023; Bai et al., 2023; Jiang et al., 2023, 2024; Cahyawijaya et al., 2024; Cohere, 2024, interalia). Multilingual evaluation is challenging due to the small number of multilingual benchmarks available, the lack of language diversity in them Ahuja et al. (2022) and the evidence of possible contamination of many of these benchmarks Ahuja et al. (2024). Additionally, many multilingual benchmarks are translations of benchmarks originally created in English, leading to loss of linguistic and cultural context.
@@
In this work, we perform 90K human evaluations - the largest scale multilingual human evaluation of LLMs as per our knowledge. We perform evaluation on a new set of general and culturally nuanced prompts created independently by native speakers for each language. We use a setting similar to the LMSys ChatbotArena Chiang et al. (2024) and ask human evaluators employed by Karya to perform two evaluation tasks: comparative evaluations between models, and individual evaluations or direct assessments of 30 models. Karya employs workers from all states of India, with a focus on rural and marginalized communities, making our study the first effort as per our knowledge that includes these communities in the evaluation process. In addition to performing human evaluations, we build upon prior work on LLMs as multilingual evaluators Hada et al. (2024b, a) to perform the same evaluations using LLMs as judges. We also use LLMs to perform safety evaluation, for which we do not engage Karya workers due to ethical concerns.
@@ 2 Related Work
Kakwani et al. (2020) release the first Indic NLU benchmark, IndicGLUE, for 11 languages. Doddapaneni et al. (2023) build on top of the former and release IndicXTREME, spanning all 22 languages. On the NLG side, Kumar et al. (2022) offer IndicNLGsuite, covering 5 tasks across 11 languages. Gala et al. (2023) release a machine translation benchmark, IN22, for both conversational and general translation evaluation across all 22 languages. Recently, Singh et al. (2024a) put forth IndicNLGBench, a collection of diverse generation tasks like cross-lingual summarization, machine translation, and cross-lingual question answering.
@@
LLMs have been shown to be useful as evaluators due to their instruction following abilities, but studies have also shown that they can be biased and may not always agree with human judgments. Hada et al. (2024b, a) conduct a comprehensive survey of LLMs as an evaluators in the multilingual setting, and also release, METAL, a benchmark for LLM-based Summarization evaluation across 10 languages. Other recent works such as Liu et al. (2024); Shen et al. (2023); Kocmi and Federmann (2023) also discuss and use LLMs for evaluations at scale, and Zheng et al. (2023) employ GPT-4 as an evaluator alongside humans to build the MTBench and ChatbotArena leaderboard. Ning et al. (2024) propose an LLM-based peer-review process to automatically evaluate the outputs of an LLM, by other models in the setup.
@@ 3.1 Prompt Curation
We include the following 10 Indian languages in our evaluation: Hindi, Tamil, Telugu, Malayalam, Kannada, Marathi, Odia, Bengali, Gujarati, and Punjabi. Our prompts comprise of 20 questions per language - 5 on health, 5 on finance, and 10 culturally nuanced prompts that were created independently by native speakers in the research team and Karya . Although we currently evaluate only on a small set of prompts, we plan to scale the number of prompts by allowing evaluators to create their own prompts similar to ChatbotArena Yang et al. (2023).
@@ 3.2 Model Selection
We evaluate popular Indic language models in addition to the leading proprietary LLMs. Most of the Indic LLMs are fine-tuned versions of the open-source Llama-2 7B base model Touvron et al. (2023), Mistral 7B Jiang et al. (2023) or Gemma 7B Mesnard and Team (2024) models, hence we added the instruct versions of these models to our evaluation to determine the gain obtained by fine-tuning these models with Indic data. We have also included the latest Llama-3 8B and Llama-3 70B AI@Meta (2024) models to evaluate their effectiveness for multilingual fine-tuning. We list all models under consideration Appendix B in Table 2 and Table 3. We are aware that it is not entirely fair to compare open-source models with API-based systems that may have several other components in place, such as language detectors, more sophisticated safety guardrails etc., however, we treat all models as the same for this study and urge the reader to keep this in mind while interpreting the results. The details for generating model query-response pairs can be found in Appendix C.
@@ 3.3 Evaluation Setup
We evaluate the generated model query-response using two different strategies and by two types of evaluators. First, we do a pairwise comparison (battle) between model responses for the same prompt and calculate Elo Ratings Elo (1978); Boubdir et al. (2023). Second, we also calculate various direct assessment metrics for each model prompt-response data point. We evaluate 12-15 models for each language except Hindi for which we evaluate 20 models. The detailed statistics of the evaluation datapoints can be seen in Appendix F.
We use human evaluation to annotate the datapoints. Each datapoint is annotated by three annotators and the majority vote is taken. If all three votes are different, we treat it as a tie in case of a battle and take average score in case of direct assessment metrics. We also use an LLM (GPT-4-32K) for annotating the battles as well as calculating the direct assessment metrics. The instructions are provided in English and a detailed description of the task and scoring rubric is also provided.
@@ 3.3.1 Pairwise comparison
The annotators perform the evaluation task on a smartphone. The annotators are provided with the query, the two model responses (model names are hidden), and set of three options - A (response 1 is better), B (response 2 is better), and C (tie, equally good/bad). We also ask the annotators to provide a spoken justification for the chosen response that is captured as audio by the app. The annotation guidelines and Hindi app screenshots are available in Appendix E.1.
@@
We also evaluate battles using GPT-4-32K as an LLM evaluator. The setting is similar to the one provided to humans. The detailed prompt is provided in Figure 10.
@@ 3.3.3 Safety Evaluation
We use the Hindi prompts from RTP-LX de Wynter et al. (2024) dataset [^3] https://github.com/microsoft/RTP-LX which is specifically designed to elicit toxic responses and ask the relevant models to generate completions. These completions are then evaluated using an LLM evaluator with the same prompt used for individual evaluations (Fig 11). The detailed rubric for Safety is defined in Fig 15. We also perform an exact match with the Hindi block words from the FLORES Toxicity-200 dataset [^4] https://github.com/facebookresearch/flores/blob/main/toxicity/README.md Costa-jussà et al. (2022) to check for toxic words in the output.
@@ 3.4 Inter-Annotator Agreement
To check for the quality of human annotation, we calculate inter-annotator agreement between the three human annotators using two metrics - Percentage Agreement (PA) and Fleiss Kappa ($\kappa$). These metrics are also used to judge the alignment between humans and LLMs for the evaluation tasks, following the same setup as prior work Hada et al. (2024b, a). We also calculate the correlation between rankings of the leaderboards obtained from human and LLM evaluations using Kendall’s Tau ($\tau$).
@@ 4.1 Leaderboard Analysis
Figure 2 depicts a visualization of the leaderboard based on the MLE Elo rating method discussed in Section 3.3.1. For the Direct Assessment scores, we report the average score across all query-response pairs for a model in Figure 3. We include both human and LLM-evaluator leaderboards in these visualizations. For the safety evaluation, the scores depict the fraction of prompts for which models gave problematic content. A detailed description of how each leaderboard is constructed along with the scores is available in Appendix G.
@@
Proprietary LLMs like GPT-4 and Gemini-Pro 1.0555available only for Hindi and Bengali rank in the upper middle portion of the human evaluations leaderboard, and top most of the LLM-evaluator leaderboards, showing evidence of self-bias by GPT-4 Panickssery et al. (2024); Xu et al. (2024). GPT-3.5-Turbo, however, ranks across the lower-middle half and performs worse than the finetuned Indic LLMs. The LLM evaluator also tends to favour Gemma 7B more than humans, suggesting that there may be some artifacts in some models that the LLM-evaluator picks up on. [^6] https://lmsys.org/blog/2024-05-08-Llama3/ We also notice that Elo ratings by humans tend to be lower than the ones given by LLM overall. This can be attributed to the fact that LLMs pick fewer ties and tend to be more decisive in comparison to humans Sharma et al. (2024); Hosking et al. (2024); Wu and Aji (2023).
@@ 4.2 RTP-LX Safety Analysis
We conduct a safety analysis of all the Hindi LLMs, and evaluate the completions using GPT-4-32K. Following Hada et al. (2024a), we use a temperature of 1.0 to elicit as lower-probability generations as possible, which might be problematic. API based LLMs, such as GPT and Gemini-Pro usually have guardrails and content moderation services before the actual model, and hence, we find that our prompts are blocked. Figure 4 shows the fraction of toxic/problematic completions for each model, as evaluated by GPT-4 and a heuristic word match from the Toxicity-200 block list.
We find that the heuristic word match fails to identify several cases of toxic completion as the the word list is limited and contains mostly stem forms of the toxic word, and other forms of the word are bypassed. GPT-3.5-Turbo produces the least toxic completions ($\sim$ 10%), followed by GPT-4o and GPT-4. The AryaBhatta-Gemma models produce the highest number of toxic completions ($\sim$ 60%), while the Aya Model generated the most toxic/profane content in its generations upon manual checking.
@@ 4.3 Agreement between LLM and Humans
Next, we analyze the agreement between humans and LLM evaluator across the two types of evaluations. We compute the Percentage Agreement (PA) and Fleiss Kappa ($\kappa$) score which are calculated at a per-datapoint level as well as the general agreement between the leaderboards using Kendall’s Tau ($\tau$). The PA score is reported in Appendix H.
Query Type Pairwise Direct
@@
Table 1: Average Fleiss Kappa ($\kappa$) correlations between Humans and Human-LLM for both evaluations across query types. Here $\mathcal{H}$ stands for Humans.
@@
On average humans have a moderate $\kappa$ score [^7] Generally, $\kappa>0.45$ is considered strong positive agreement of 0.54 whereas the human-average and LLM have a $\kappa$ score of 0.49. An ablation on the query-type in Table 1 reveals that LLM evaluator agrees comparatively less on the culturally-nuanced queries. A language-wise breakdown of the $\kappa$ scores can be seen in Figure 5. For pairwise evaluation, humans tend to have higher agreements among themselves than with the LLM across all languages except for Hindi and Kannada. The LLM evaluator has very low agreement with humans on Marathi, Bengali and Punjabi.
@@
In this case, humans tend to have a slightly lower but similar agreement to the pairwise scores. However, the agreement between humans and LLMs significantly drops and is the lowest again for the culturally-nuanced set of queries. Figure 5 shows that for Direct Assessment, humans have similar agreement across all languages but the human-LLM agreement is significantly lower particularly for Bengali and Odia. This indicates that Direct Assessment may be a harder evaluation task for LLMs.
@@
We check the agreement between the leaderboards to get a sense of agreement on higher level trends. We report the Kendall Tau ($\tau$) scores between the rankings of models in both Elo and Direct Assessment leaderboards in Table 36. On average we see a high $\tau$ score [^8] Generally, $\tau>0.7$ is considered a strong positive correlation of 0.76 for the Elo leaderboards which signifies that the human and LLM-evaluator agree on the general trends. From Figure 2 we can see that although the absolute rankings are not same, we can still find similar sub-group of models. We see this agreement go down for the Direct Assessment leaderboard which has an average agreement of 0.65. This reinforces our hypothesis that the LLM evaluator is worse at the Direct Assessment task.

### v2

# Pariksha: Large-Scale Investigation of Human-LLM Evaluator Agreement on Multilingual and Multi-Cultural Data

@@ Abstract
Evaluation of multilingual Large Language Models (LLMs) is challenging due to a variety of factors -- the lack of benchmarks with sufficient linguistic diversity, contamination of popular benchmarks into LLM pre-training data and the lack of local, cultural nuances in translated benchmarks. In this work, we study human and LLM-based evaluation in a multilingual, multi-cultural setting. We evaluate 30 models across 10 Indic languages by conducting 90K human evaluations and 30K LLM-based evaluations and find that models such as GPT-4o and Llama-3 70B consistently perform best for most Indic languages. We build leaderboards for two evaluation settings - pairwise comparison and direct assessment and analyse the agreement between humans and LLMs. We find that humans and LLMs agree fairly well in the pairwise setting but the agreement drops for direct assessment evaluation especially for languages such as Bengali and Odia. We also check for various biases in human and LLM-based evaluation and find evidence of self-bias in the GPT-based evaluator. Our work presents a significant step towards scaling up multilingual evaluation of LLMs.111 https://aka.ms/pariksha
@@
Pariksha: Large-Scale Investigation of Human-LLM Evaluator Agreement on Multilingual and Multi-Cultural Data
@@ 1 Introduction
Figure 1: Evaluation pipeline: (1) We curate a diverse set of evaluation prompts with the help of native speakers. (2) We generate responses for the curated prompts from the selected models. (3) We evaluate generated responses in two settings (direct assessment and pairwise comparison) by both humans and an LLM. (4) We construct leaderboards using scores obtained and analyze the agreement between human and LLM evaluators.
@@
However, standard benchmarking suffers from the following issues: many popular benchmarks are available on the web and have already been consumed in the training data of LLMs, rendering them unsuitable for fair evaluation. This phenomenon is known as test dataset contamination, and recent works Ravaut et al. (2024); Golchin and Surdeanu (2024); Dong et al. (2024); Oren et al. (2024); Deng et al. (2024) have suggested that contamination may occur not only during pre-training, but also during fine-tuning and evaluation Balloccu et al. (2024). This calls for dynamic benchmarking with the help of humans Chiang et al. (2024); Yang et al. (2023). Although, human evaluation is considered the gold standard, it can be expensive and time consuming. Due to this, the use of LLM-evaluators, where an LLM itself is used to evaluate the output of another LLM (sometimes the same LLM) has become very popular.
@@
Most studies on LLM training and evaluation focus on English. Recent works have shown that LLMs perform worse on non-English languages, particularly those written in scripts other than the Latin script, and under-resourced languages Ahuja et al. (2023, 2024); Asai et al. (2024). Studies on cultural values in LLMs have also shown that frontier models such as GPT-4 align more closely to Western, Rich, and Industrialized norms Rao et al. (2023). This has led to a proliferation of models being built for specific languages, cultures and regions such as Indic, Arabic, African, Chinese, European, and Indonesian (Gala et al., 2024; Sengupta et al., 2023; Zeng et al., 2023; Bai et al., 2023; Cahyawijaya et al., 2024; Cohere, 2024; Üstün et al., 2024, interalia). Multilingual evaluation is challenging due to the small number of multilingual benchmarks available, the lack of language diversity in them Ahuja et al. (2022) and the evidence of possible contamination of many of these benchmarks Ahuja et al. (2024). Additionally, many multilingual benchmarks are translations of benchmarks originally created in English, leading to loss of linguistic and cultural context.
@@
In this work, we perform 90K human evaluations - the largest scale multilingual human evaluation of LLMs as per our knowledge. We perform evaluation on a new set of general and culturally-nuanced prompts created independently by native speakers for each language. We use a setting similar to the LMSys ChatbotArena Chiang et al. (2024) and ask human evaluators employed by Karya 222 https://karya.in , an ethical data company, to perform two evaluation tasks: comparative evaluations between models, and individual evaluations or direct assessments of 30 models. Karya employs workers from all states of India, with a focus on rural and marginalized communities, making our study the first effort as per our knowledge that includes these communities in the evaluation process. In addition to performing human evaluations, we build upon prior work on LLMs as multilingual evaluators Hada et al. (2024b, a) to perform the same evaluations using LLMs as judges. We also use LLMs to perform a preliminary safety evaluation, for which we do not engage Karya workers due to ethical concerns.
@@ 2 Related Work
Kakwani et al. (2020) release the first Indic NLU benchmark, IndicGLUE, for 11 languages. Doddapaneni et al. (2023) build on top of the former and release IndicXTREME, spanning all 22 languages. On the NLG side, Kumar et al. (2022) offer IndicNLGsuite, covering 5 tasks across 11 languages. Gala et al. (2023) release a machine translation benchmark, IN22, for both conversational and general translation evaluation across all 22 languages. Recently, Singh et al. (2024a) put forth IndicGenBench, a collection of diverse generation tasks like cross-lingual summarization, machine translation, and cross-lingual question answering.
@@
LLMs have been shown to be useful as evaluators due to their instruction following abilities, but studies have also shown that they can be biased and may not always agree with human judgments. In prior work (Hada et al. (2024b, a)), we conducted a comprehensive survey of LLMs as an evaluators in the multilingual setting, and also released METAL, a benchmark for LLM-based Summarization evaluation across 10 languages. Other recent works such as Liu et al. (2024); Shen et al. (2023); Kocmi and Federmann (2023) also discuss and use LLMs for evaluations at scale, and Zheng et al. (2023) employ GPT-4 as an evaluator alongside humans to build the MTBench and ChatbotArena leaderboard. Ning et al. (2024) propose an LLM-based peer-review process to automatically evaluate the outputs of an LLM, by other models in the setup.
@@ 3.1 Prompt Curation
India is a diverse country where the language and cultural nuances change every few kilometers. This diversity necessitates involving native speakers to create and evaluate these prompts. We include the following 10 Indian languages in our evaluation: Hindi, Tamil, Telugu, Malayalam, Kannada, Marathi, Odia, Bengali, Gujarati, and Punjabi. Our prompts comprise of 20 questions per language - 5 on health, 5 on finance, and 10 culturally-nuanced prompts - that were created by native speakers in the research team and Karya  (Table 1). Although we currently evaluate only on a small set of prompts, we plan to scale the number of prompts by allowing evaluators to create their own prompts similar to ChatbotArena Yang et al. (2023).
The prompts were created independently for each language following the same guidelines and are not translations. While the finance and health prompts are similar across languages, cultural prompts may differ slightly due to unique nuances of each language. This differentiation is crucial to capture the distinct aspect of each language and evaluate Indic LLMs in context.
Prompt Type Examples
Finance (5) What is the difference between a debit card and a credit card?
Health (5) How can I improve my posture to prevent back and neck pain?
Cultural (10) (Kannada) Although in the neighboring states movie actors rise to prominence in politics, why is it not seen in Karnataka?
(Telugu) In Telugu tradition, why do parents of girls pay for the first birth of a child?
Table 1: Table containing number of prompts in each category per language with examples (English translation for readability).
@@ 3.2 Model Selection
We evaluate popular Indic language models in addition to the leading proprietary LLMs. Most of the Indic LLMs are fine-tuned versions of the open-source Llama-2 7B base model Touvron et al. (2023), Mistral 7B Jiang et al. (2023) or Gemma 7B Mesnard and Team (2024) models, hence we added the instruct versions of these models to our evaluation to determine the gain obtained by fine-tuning these models with Indic data. We have also included the latest Llama-3 8B and Llama-3 70B AI@Meta (2024) models to evaluate their effectiveness for multilingual fine-tuning. We list all models under consideration in Appendix B in Table 5 and Table 6.
We are aware that it is not entirely fair to compare open-source models with API-based systems that may have several other components in place, such as language detectors, more sophisticated safety guardrails etc., however, we treat all models as the same for this study and urge the reader to keep this in mind while interpreting the results. The details for generating model query-response pairs can be found in Appendix C.
@@ 3.3 Evaluation Setup
We evaluate the models on open-ended Question Answering on the aforementioned prompts using two different strategies and by two types of evaluators. First, we do a pairwise comparison (battle) between model responses for the same prompt and calculate Elo Ratings Elo (1978); Boubdir et al. (2023). Second, we also calculate various direct assessment metrics for each model prompt-response data point. A total of 21690 datapoints (battles) are evaluated by LLMs evaluator for pairwise evaluation whereas 2880 model query-response pairs are evaluated for 3 metrics (hallucinations, task quality and linguistic acceptability) which results to 8640 datapoints. Hence, a total of 21690 + 8640 = 30330 datapoints. We evaluate 12-15 models for each language except Hindi for which we evaluate 20 models. The detailed statistics of the evaluation datapoints can be seen in Table 2.
Each datapoint is evaluated by three human annotators and the majority vote is taken, yielding an overall of 3 $\times$ 30330 = 90K annotations. If all three votes are different, we treat it as a tie in case of a battle and take average score in case of direct assessment metrics. In addition to human evaluation, we also use an LLM (GPT-4-32K) for evaluating the battles as well as providing scores using the direct assessment metrics.
@@ 3.3.1 Pairwise comparison
The annotators perform the evaluation task on a smartphone. The annotators are provided with the query, the two model responses (model names are hidden), and set of three options - A (response 1 is better), B (response 2 is better), and C (tie, equally good/bad). During evaluation, we ask the annotators to provide a spoken justification for the chosen response that is captured as audio by the app. The annotation guidelines and Hindi app screenshots are available in Appendix E.1.
@@
We also evaluate battles using GPT-4-32K as an LLM evaluator. The setting is similar to the one provided to humans with the same rubric and a similar set of instructions provided as a prompt to the LLM. The detailed prompt is provided in Figure 10.
@@ 3.3.3 Safety Evaluation
We also conduct a preliminary safety evaluation to estimate the efficacy of guardrails in different LLMs. For this, we use the Hindi prompts from RTP-LX [^4] https://github.com/microsoft/RTP-LX de Wynter et al. (2024) which is specifically designed to elicit toxic responses and ask the models under consideration to generate completions. These completions are then evaluated using an LLM evaluator with the same prompt used for individual evaluations (Figure 11). We do not employ humans for the safety evaluation due to ethical concerns. The detailed rubric for Safety is defined in Fig 15. We also perform an exact match with the Hindi block words from the FLORES Toxicity-200555https://github.com/facebookresearch/flores/blob/main/toxicity/README.md Costa-jussà et al. (2022) to check for toxic words in the output.
@@ 3.4 Inter-Annotator Agreement
To check for the quality of human annotation, we calculate inter-annotator agreement between the three human annotators using two metrics - Percentage Agreement (PA) and Fleiss Kappa ($\kappa$). These metrics are also used to judge the alignment between humans and LLMs for the evaluation tasks, following the same setup as our prior work Hada et al. (2024b, a). We also calculate the correlation between rankings of the leaderboards obtained from human and LLM evaluations using Kendall’s Tau ($\tau$).
@@ 4.1 Leaderboard Analysis
Figure 2 depicts a visualization of the leaderboard based on the MLE Elo rating method discussed in Section 3.3.1. For the Direct Assessment scores, we report the average score across all query-response pairs for a model in Figure 3. We include both human and LLM-evaluator leaderboards in these visualizations. For the safety evaluation, the scores depict the fraction of prompts for which models gave problematic content. A detailed description of how each leaderboard is constructed along with the scores is available in Appendix F.
@@
Proprietary LLMs like GPT-4 and Gemini-Pro 1.0666available only for Hindi and Bengali rank in the upper middle portion of the human evaluations leaderboard, and top most of the LLM-evaluator leaderboards, showing evidence of self-bias by GPT-4 Panickssery et al. (2024); Xu et al. (2024). GPT-3.5-Turbo, however, ranks across the lower-middle half and performs worse than the fine-tuned Indic LLMs. The LLM evaluator also tends to favour Gemma 7B more than humans, suggesting that there may be some artifacts in some models that the LLM-evaluator picks up on. [^7] https://lmsys.org/blog/2024-05-08-Llama3/ We also notice that Elo ratings by humans tend to be lower than the ones given by LLM overall. This can be attributed to the fact that LLMs pick fewer ties and tend to be more decisive in comparison to humans Sharma et al. (2024); Hosking et al. (2024); Wu and Aji (2023).
@@ 4.2 RTP-LX Safety Analysis
We present the preliminary safety analysis of all the Hindi LLMs, for which we evaluate the completions using GPT-4-32K. Following Hada et al. (2024a), we use a temperature of 1.0 to elicit even unexpected generations which might be problematic. API based LLMs, such as GPT and Gemini-Pro usually have guardrails and content moderation services before the actual model, and hence, we find that our prompts are blocked. Figure 4 shows the fraction of toxic/problematic completions for each model, as evaluated by GPT-4 and a heuristic word match from the Toxicity-200 block list.
We find that the heuristic word match fails to identify several cases of toxic completion as the the word list is limited and contains mostly stem forms of the toxic word, and other forms of the word are bypassed. GPT-3.5-Turbo produces the least toxic completions ($\sim$ 10%), followed by GPT-4o and GPT-4. We also note that the Gemma model and its fine tuned variants (Aryabhatta variants and Navarasa) consistently perform the worst. However, since these evaluations are automated, they may contain potential biases. We leave further study with human evaluations as part of future work.
@@ 4.3 Agreement between LLM and Humans
Next, we analyze the agreement between humans and LLM evaluator across the two types of evaluations. We compute the Percentage Agreement (PA) and Fleiss Kappa ($\kappa$) score which are calculated at a per-datapoint level as well as the general agreement between the leaderboards using Kendall’s Tau ($\tau$). The PA score is reported in Appendix G.
Prompt Type Pairwise Direct
@@
Table 3: Average Fleiss Kappa ($\kappa$) correlations between Humans and Human-LLM for both evaluations across prompt types. Here $\mathcal{H}$ stands for Humans.
@@
On average humans have a moderate $\kappa$ score [^8] Generally, $\kappa>0.45$ is considered strong positive agreement of 0.54 whereas the human-average and LLM have a $\kappa$ score of 0.49. An ablation on the prompt-type in Table 3 reveals that LLM evaluator agrees comparatively less on the culturally-nuanced prompt. A language-wise breakdown of the $\kappa$ scores can be seen in Figure 5. For pairwise evaluation, humans tend to have higher agreements among themselves than with the LLM across all languages except for Hindi and Kannada. The LLM evaluator has very low agreement with humans on Marathi, Bengali and Punjabi.
@@
In this case, humans tend to have a slightly lower but similar agreement to the pairwise scores. However, the agreement between humans and LLMs significantly drops and is the lowest again for the culturally-nuanced set of prompts. Figure 5 shows that for Direct Assessment, humans have similar agreement across all languages but the human-LLM agreement is significantly lower particularly for Bengali and Odia. This indicates that Direct Assessment may be a harder evaluation task for LLMs.
@@
We check the agreement between the leaderboards to get a sense of agreement on higher level trends. We report the Kendall Tau ($\tau$) scores between the rankings of models in both Pairwise (Elo) and Direct Assessment (DA) leaderboards in Table 4. On average we see a high $\tau$ score [^9] Generally, $\tau>0.7$ is considered a strong positive correlation of 0.76 for the Elo leaderboards which signifies that the human and LLM-evaluator agree on the general trends. From Figure 2 also we can see that although the absolute rankings are not same, we can still find similar sub-group of models. We see this agreement go down for the DA leaderboard which has an average agreement of 0.65. We also report a detailed language-wise breakdown and find that for Elo leaderboards, Gujarati has the highest correlation while Bengali is the lowest. For the DA leaderboards, the scores drop and it is again the lowest for Bengali. This reinforces our hypothesis that the LLM evaluator is worse at the DA task.